Lighting manufactures a variety of lamps for lighting specialty stores. In​ September, Parker Lighting received an order from Zorbet Decor for 60 ​mission-style table lamps​ (Model 5A). The order from Zorbet Decor became Job Number 804 at Parker Lighting.

Materials requisition​ #1247 for Job 804 shows that a total of $309 of direct materials were used in manufacturing the job. In addition to the materials​ requisition, the labor time records show that a total of $105 of direct labor costs was incurred in producing these​ lamps; a total of 15 direct labor hours were worked on this job.

Requirements: Add direct materials and direct labor to the Job Cost Record for Job 804. Manufacturing overhead has already been added to the Job Cost Record for this job. Assume that the job was completed on September 19. Complete the Job Cost Record by calculating the total job cost and the cost per unit. Remember that this job consisted of 60 lamps​ (units).

Requirement Number: 1247

Job Cost Record

Job Number:

804

Customer:

Zorbet Decor

Job description:

60 Model 5A - Mission-style table lamps

Date Started:

Sep. 14

Date Completed:

Sep. 19

Manufacturing Cost Information:

Cost Summary

Direct Materials

Req. # :

1247

Direct Labor

Manufacturing Overhead

15 hours × $4 per direct labor hour

$60

Total Job Cost

Number of Units

÷

Cost per Unit
Answer #1

Calculation of the total job cost= Materials Cost + direct labour + Manufacturing Overhead

= $309 + $105 + ($60*6 Days) = 309+105+360

= $774

Calculation of cost per unit = Total cost / No of units

=  774/60

= $ 12.9 per unit